# Native Design Bias: Studying the Impact of English Nativeness on Language Model Performance

### v2

@@ 5 Results
In line with our hypotheses, the results indicate performance differences between native and non-native English speakers and Western and not Western native speakers for GPT4o, Sonnet, and Qwen for the objective classification tasks, as shown in Table 1. However, the generative tasks indicate an opposite preference as displayed in Table 4. Informing the models about the user’s nativeness further reduces answer quality for non-native English speakers, also resulting in a strong anchoring effect.
5.1 Classification tasks
Do LLMs perform better for certain groups on both objective and subjective classification tasks? Table 1 shows the average performance and standard deviation over three runs on the objective and subjective classification tasks for the different models and groups. The subjective classification task entails all tasks without an objectively correct answer, as this depends on subjective interpretation. The only subjective task included in our paper is the Amazon Food review task.
objective tasks subjective tasks
native non-native western native not western native native non-native western native not western native
GPT3.5 $0.82\pm 0.01$ $0.81\pm 0.01$ $0.81\pm 0.00$ $0.81\pm 0.01$ $0.76\pm 0.01$ $0.80\pm 0.01$ $0.76\pm 0.00$ $0.80\pm 0.00$
GPT4o $0.94\pm 0.00$ $0.93\pm 0.00$ $0.95\pm 0.01$ $0.93\pm 0.01$ $0.77\pm 0.00$ $0.84\pm 0.01$ $0.77\pm 0.01$ $0.83\pm 0.01$
Haiku $0.87\pm 0.01$ $0.86\pm 0.00$ $0.86\pm 0.00$ $0.87\pm 0.00$ $0.75\pm 0.00$ $0.80\pm 0.00$ $0.71\pm 0.00$ $0.81\pm 0.00$
Sonnet $0.87\pm 0.00$ $0.87\pm 0.00$ $0.87\pm 0.01$ $0.87\pm 0.00$ $0.77\pm 0.00$ $0.79\pm 0.01$ $0.76\pm 0.01$ $0.80\pm 0.00$
Qwen 7B $0.83\pm 0.00$ $0.82\pm 0.01$ $0.83\pm 0.00$ $0.82\pm 0.00$ $0.75\pm 0.00$ $0.79\pm 0.00$ $0.72\pm 0.00$ $0.79\pm 0.00$
Table 1: Overall average accuracy of the objective and subjective classification tasks per model and group averaged over three runs and including the standard deviation over these different runs. Significant differences with $p<0.05$ for a Wilcoxon signed-rank test are indicated in bold.
We find that for the objective classification tasks, only GPT4o, Sonnet, and Qwen 7B perform significantly better for the native group of annotators. For the other models, no significant differences are found between the groups. For the western native group, we find that only GPT4o and Qwen 7B a significantly higher performance for the Western native English speakers. Interestingly, we thus conclude that Sonnet is less tailored towards Western native English, which suggests efforts to take into account different English dialects. GPT4o on the other hand, is tailored towards native English speakers, and more specifically to Western native speakers. Interestingly, Qwen, which is predominantly trained on Chinese text, also prefers native and Western native English speakers.
For both the Claude and GPT models, we find that the larger model of the model family is more sensitive to the subtle differences in the prompts of native and non-native speakers. In literature, we find similar positive correlation between model size and bias that slips into models in terms of gender bias Tal et al. (2022). Moreover, Sclar et al. (2023) even show that prompt sensitivity does not decrease for larger model size.
Surprisingly, Table 1 shows the opposite effect for the subjective classification tasks, preferring both the non-native and not Western native English speakers. This finding is remarkable, given the results in the subjective classification literature (Santy et al., 2023; Durmus et al., 2023).
@@ 5.1 Classification tasks
To understand this contrasting phenomenon, we first analyze the results for both the native and non-native English speakers, and the Western native and not Western native English speakers, in depth. We find that for the native English-speaking group, the models often predict the rating more positively than actually intended, and this effect is even more pronounced for the Western native English speakers. While for the non-native English group and not Western native English speaking group, GPT4o predicted 50% of all wrongly predicted annotations to be more positive than intended, this was around 60% for the native English speaking group and even 70% for the Western native English speaking group for GPT4o indicating cultural differences. Appendix H includes the different distributions per model.
What is the effect on the performance when informing the model about the nativeness of the annotator? For these experiments, we added the user’s nativeness to the system prompt. More specifically, we asked the model to respond as if interacting with a (non-)native English speaker.
objective tasks subjective tasks
correct information wrong information correct information wrong information
native non-native native non-native native non-native native non-native
GPT3.5 $0.84\pm 0.00$ $0.81\pm 0.01$ $0.83\pm 0.01$ $0.83\pm 0.01$ $0.76\pm 0.00$ $0.83\pm 0.01$ $0.78\pm 0.00$ $0.83\pm 0.00$
GPT4o $0.94\pm 0.00$ $0.93\pm 0.00$ $0.94\pm 0.00$ $0.93\pm 0.00$ $0.77\pm 0.01$ $0.82\pm 0.01$ $0.77\pm 0.01$ $0.84\pm 0.00$
Haiku $0.87\pm 0.01$ $0.84\pm 0.00$ $0.85\pm 0.00$ $0.86\pm 0.00$ $0.75\pm 0.00$ $0.80\pm 0.00$ $0.74\pm 0.00$ $0.81\pm 0.01$
Sonnet $0.88\pm 0.00$ $0.63\pm 0.00$ $0.63\pm 0.00$ $0.87\pm 0.00$ $0.77\pm 0.00$ $0.78\pm 0.00$ $0.78\pm 0.00$ $0.80\pm 0.00$
Qwen 7B $0.83\pm 0.00$ $0.82\pm 0.00$ $0.83\pm 0.00$ $0.83\pm 0.00$ $0.75\pm 0.00$ $0.79\pm 0.00$ $0.75\pm 0.00$ $0.79\pm 0.00$
Table 2: Overall average accuracy and standard deviation of the objective and subjective classification tasks per model and group averaged over three runs when adding information about the annotator’s nativeness to the prompt. Significant differences with $p<0.05$ for a Wilcoxon signed-rank test are indicated in bold.
Table 2 shows the effect of adding (in)correct information about the nativeness of the annotator to the system prompt on the performance of the models. Compared to Table 1, we now find that all models significantly perform better for the objective classification tasks for the native group compared to the non-native group when correct information is added. When inaccurately labeling native speakers as non-native and vice versa, we see that only GP4o and Qwen still significantly perform better for the native group, suggesting that they are less affected by this additional information. Interestingly, Sonnet is most affected by the information about the (in)correct nativeness of the prompt writer. We find that Sonnet answered several questions in languages other than English, such as Spanish, French or Indonesian, when responding as if interacting with non-native speakers. This is remarkable and shows a clear anchoring effect, considering we only instructed the model to respond as if interacting with a non-native English speaker. Interestingly, this did not occur for the other models
Moreover, also for the other two models that are mostly affected by this anchoring effect, we manually analyzed the errors. GPT3.5 makes more mistakes when informed about the prompt writer being non-native, due to repetition of the instructions, rather than answering the question. Haiku explains the answers, arguing why one option is better than another, thereby failing to follow the instructions, making it hard to assess performance automatically. If both answers are mentioned, we classify the response as inaccurate. Note that this anchoring effect is only seen for the objective classification tasks. It only slightly impacts the subjective classification tasks and does not result in new significant differences than those shown in Table 1. This overall anchoring effect, shows that these three models (GPT3.5, Haiku, and Sonnet) are still sensitive to (useless) information that is given within the prompt as also shown by  Shi et al. (2023); Jones and Steinhardt (2022)
What is the result of first asking the model to guess about the nativeness of the prompt writer? In these experiments, the model first guesses the user’s nativeness from their annotation, and then generates a task response within the same chat.
objective tasks subjective tasks
guess native guess non-native guess native correctly guess non-native correctly guess native guess non-native guess native correctly guess non-native correctly
GPT3.5 $0.80\pm 0.01$ $0.77\pm 0.00$ $0.75\pm 0.02$ $0.72\pm 0.01$ $0.75\pm 0.01$ $0.75\pm 0.01$ $0.77\pm 0.03$ $0.77\pm 0.02$
GPT4o $0.85\pm 0.00$ $0.84\pm 0.01$ $0.82\pm 0.05$ $0.82\pm 0.02$ $0.77\pm 0.01$ $0.83\pm 0.01$ $0.74\pm 0.1$ $0.71\pm 0.02$
Haiku $0.65\pm 0.00$ $0.64\pm 0.01$ $0.47\pm 0.03$ $0.66\pm 0.01$ $0.67\pm 0.00$ $0.66\pm 0.01$ $0.74\pm 0.05$ $0.51\pm 0.04$
Sonnet $0.63\pm 0.01$ $0.62\pm 0.01$ $0.52\pm 0.01$ $0.51\pm 0.01$ $0.72\pm 0.01$ $0.77\pm 0.00$ $0.64\pm 0.12$ $0.75\pm 0.04$
Qwen 7B $0.69\pm 0.00$ $0.69\pm 0.00$ $0.74\pm 0.00$ $0.74\pm 0.00$ $0.67\pm 0.00$ $0.71\pm 0.00$ $0.44\pm 0.00$ $0.77\pm 0.00$
Table 3: Overall average accuracy and standard deviation over the three runs of all classification tasks per model and group when guessed by the model whether the person is a native or non-native speaker, and when it is guessed correctly. Significant differences with $p<0.05$ for a Wilcoxon signed-rank test are indicated in bold. No statistical testing is done for the correctly guessed category for the subjective tasks due to the low amount of observations.
Table 3 shows that for GPT3.5, GPT4o, and Sonnet, first explicitly letting the model guess about the nativeness of the prompt writer, results in a significantly better score for the prompts from native than from non-native English speakers, regardless of having correctly guessed the nativeness. Compared to our initial experiments, we thus see that GPT3.5 is more led by this added question and its information and starts performing better for the native class. We find that both GPT models provide accurate explanations of the models’ guesses, recognizing words within the prompt as i.e. British English. Generally, non-native prompts are identified through grammatical errors or awkward phrasing.
The performance difference for Haiku between the two correctly guessed groups is remarkable. The model has predicted many well-performing non-native annotations to be actually native for the objective classification tasks. Qwen now shows a non-significant difference between both. We find that these models struggle with following the guideline to respond with only (non-)native explaining why one is more likely than the other, as does Sonnet. Additionally, Qwen often refrains from guessing, stating it is difficult to determine the nativeness. This results in a lower performance and a non-significant difference between both groups. For the subjective tasks, we find that GPT4o, Sonnet, and Qwen are robust to this prompt, still providing significant differences for the non-native group, while Haiku and GPT3.5 are more impacted.
5.2 Generative tasks
Do we find similar performance differences for the generative tasks?
native non-native western native not western native
GPT3.5 $0.44\pm 0.00$ $0.44\pm 0.00$ $0.44\pm 0.00$ $0.44\pm 0.00$
GPT4o $0.47\pm 0.00$ $0.47\pm 0.00$ $0.47\pm 0.00$ $0.47\pm 0.00$
Haiku $0.45\pm 0.00$ $0.46\pm 0.01$ $0.45\pm 0.01$ $0.46\pm 0.01$
Sonnet $0.45\pm 0.00$ $0.43\pm 0.02$ $0.45\pm 0.00$ $0.45\pm 0.00$
Qwen 7B $0.41\pm 0.00$ $0.44\pm 0.02$ $0.41\pm 0.01$ $0.42\pm 0.01$
Table 4: Overall average BLEU score of the generative task per model and group averaged over three runs and including the standard deviation over these different runs. Significant differences with $p<0.05$ for a Wilcoxon signed-rank test are indicated in bold.
Table 4 shows the average BLEU scores and standard deviation over the three different runs per model and group. The statistical analysis includes all performance metrics. These metrics are included in Appendix L. For the generative tasks, all models perform overall significantly better for the non-native group, and all models except for GPT4o and Sonnet prefer the not western native group.
Despite this clear preference for the non-native group, we do find that the performance highly depends on the specific task at hand. In general, the generative tasks can be divided into two groups: group one with CODA19, which includes a medical research article generation task, and the paraphrasing dataset, group two with StoryCloze and CNN Dailymail. We find that all models prefer the non-native English speakers or not Western native English speakers for group one. Most models prefer (Western) native English speakers for group two. Detailed results are shown in Table 19 in Appendix.
In interpreting these results, it is important to consider the specific nature of the generation tasks involved. CODA19 comprises medical articles that utilize specialized medical terminology. Given that most annotators were unfamiliar with this vocabulary, native English speakers did not have a specific advantage over non-native speakers. Additionally, research articles are commonly written in English by authors from various backgrounds. Therefore, it might be that this specific task is robust against the native versus non-native preference. Finally, for the paraphrasing task, through manual error analysis, we found that native and non-native English annotators handled this task differently. While native English speakers often paraphrased the sentences more freely, non-native English speakers closely followed the original structure, mainly providing synonyms and maintaining the original sentence’s format. An example is added in Appendix J.
How does informing the model about the user’s nativeness impact performance?
(non-)native system prompt guess native system prompt
correct information wrong information overall guess guess correctly
model native non-native native non-native native non-native native non-native
GPT3.5 $0.42\pm 0.01$ $0.40\pm 0.01$ $0.40\pm 0.00$ $0.42\pm 0.01$ $0.42\pm 0.00$ $0.42\pm 0.00$ $0.45\pm 0.01$ $0.42\pm 0.00$
GPT4o $0.47\pm 0.01$ $0.45\pm 0.01$ $0.45\pm 0.00$ $0.47\pm 0.00$ $0.46\pm 0.00$ $0.46\pm 0.00$ $0.48\pm 0.01$ $0.45\pm 0.01$
Haiku $0.45\pm 0.01$ $0.45\pm 0.01$ $0.45\pm 0.01$ $0.45\pm 0.01$ $0.44\pm 0.00$ $0.45\pm 0.00$ $0.34\pm 0.03$ $0.43\pm 0.00$
Sonnet $0.45\pm 0.00$ $0.45\pm 0.00$ $0.45\pm 0.00$ $0.45\pm 0.00$ $0.44\pm 0.01$ $0.45\pm 0.00$ $0.38\pm 0.02$ $0.46\pm 0.01$
Qwen 7B $0.41\pm 0.00$ $0.42\pm 0.00$ $0.41\pm 0.01$ $0.42\pm 0.01$ $0.39\pm 0.01$ $0.39\pm 0.00$ $0.43\pm 0.02$ $0.32\pm 0.01$
Table 5: Overall average BLEU score of generative tasks per model when adding correct/wrong information about whether or not the prompt was written by a native English speaker and when (correctly)guessed by the model whether the person is a native or non-native speaker. We indicated significant differences ($p<0.05$) in bold.
Interestingly, Table 5 shows that adding the correct information about the user’s nativeness affects the GPT-models largely. Now, no significant difference is found between both groups, whereas previously both models significantly preferred the non-native group. However, the result again depends on the specific task. For both GPT3.5 and GPT4o, a better performance is found for native English speakers for StoryCloze and CNN Dailymail, while the other datasets are preferred for non-native speakers. However, this difference between the native and non-native scores has increased for GPT3.5, again showing the anchoring effect. Similar to the classification tasks, this is less visible for GPT4o. While Haiku and Sonnet follow this preference for the native group on these datasets, Qwen always prefers the native group.
Additionally, when incorrect information about the user’s nativeness is introduced, all models show improved or on-par performance for the non-native group, which is presented to the model as native speakers, compared to when they are represented as non-native speakers. This is a remarkable finding, as it emphasizes the importance of the prompt over the user’s nativeness, unlike the results for the subjective classification tasks. The GPT-models are most affected by this as shown in the table.
What is the result of first asking the model to guess about the nativeness of the prompt writer?
Table 5 includes the average BLEU scores when the model (correctly) guessed the nativeness of the prompt writer. Overall, no significant differences are found for GPT3.5 and Qwen, contrary to the original experiments in Table 4. Similar to the classification tasks, we find that this is probably due to Qwen’s indecisiveness, which adds extra noise and shows the model sensitivity to useless information as in Shi et al. (2023). When looking at the guesses that were made correctly, GPT4o, Haiku, and Qwen perform significantly better for the native group than for the non-native groups. Sonnet, however still prefers the non-native group, and for GPT3.5 no significant differences are found.
@@ 6 Discussion
In our experiments, we define native bias, or western native bias, as the model’s performance disparity when prompted by native versus non-native English speakers, or Western versus not Western natives respectively. We find that there are performance differences when the model is prompted by people from different backgrounds.
We find that GPT4o, Sonnet, and Qwen perform significantly better for the native groups for the objective classification tasks. For the subjective classification task, however, we find an opposite effect, where they interpret prompts of (Western) native English speakers more positively. Similarly, the models significantly prefer the non-native group for generative tasks, but the bias varies per dataset.
Moreover, a strong anchoring effect occurs when the model knows whether or not the prompt writer is a native English speaker, either by correctly guessing, or by inserted information in the system prompt. The bias is so deeply engraved that informing the models of the wrong native and non-native groups, results in a preference towards the group that was indicated as native. The model is largely led by this added information, more than by the prompt itself. However, we find differences between the models. GPT4o appears most resistant to this anchoring effect, while Sonnet on the other hand even changes the language of the response based on this anchor. Furthermore, this anchoring effect seems to be less present for the generative tasks than for the classification tasks. For Qwen, this anchoring effect is a result of the model’s sensitivity to useless information, as this noise results in worse predictions for both groups.
Interestingly, the results also suggest that the more advanced versions of models, such as GPT4o compared to GPT3.5 and Sonnet compared to Haiku, tend to be more sensitive to the subtleties of native English. On the other hand, we also notice that GPT4o is less sensitive to added explicit information of the nativeness or non-nativeness of the annotator, and thus seems to be better hedged against explicitly added bias.

### v3

@@ 5 Results
Below, we analyze the results from our experiments answering each of the research questions. Throughout the next paragraphs, we analyze the performance of the native speakers— consisting of Western native speakers (WN) and native speakers that are non-Western (NWN)— and non-native English speakers (NN).
Figure 3: WN is best-performing for objective classification tasks and worst-performing for the subjective classification tasks. The figure shows average model performance per group and task type averaged for all models and runs; y-axis is adjusted to 0.65–1 for clarity.
The WN group performs best for the objective classification tasks, outperforming both NWN and NN.
This is shown in Figure 3, where the average performance per group on the objective classification tasks is displayed on the left. WN speakers achieve the highest overall performance in objective classification tasks, reinforcing findings from previous research hofmann2024ai; ryan2024unintended that models favor Western privileged dialects. In contrast, NWN and NN English speakers perform similarly, with the NN group slightly outperforming NWN speakers. However, this difference is minimal and not substantial enough to draw strong conclusions. The performance gap between WN and the other groups, however, suggests the advantage of Western dialects. Manual analysis reveals how LLM misclassifications stem from ambiguities in non-native prompts. In Timetravel, less fluent phrasing made incorrect options appear plausible, while in McTaco and TweetQA, non-native formulations led to misinterpretations. This highlights an inherent model bias toward native speakers rather than annotation inconsistencies.
The WN group performs worst for the subjective classification task as models predict their rating more positively than actually intended.
@@
The right part of Figure 3 shows this opposite effect for the subjective classification tasks. For these tasks, both the NN and NWN show again similar performance and are now outperforming the WN group. This finding is remarkable, as it contradicts the results in the subjective classification literature (santy-etal-2023-nlpositionality; durmus2023towards). When further analyzing the results, we find that for the Western native English-speaking group, we find that the models often predict the rating more positively than actually intended. While for the NN and NWN groups, GPT4o predicted around 50% of all wrongly predicted annotations to be more positive than intended, this was around 70% for the WN English-speaking group for GPT4o indicating cultural differences. Appendix H includes more information on the different answer distributions per model.
Figure 4: The generative tasks are more robust against native bias. This figure shows the average model performance for the generative tasks per group and metric averaged over the different runs. We rescaled the results so that they range from 0 to 1.
The generation tasks are more robust against (Western) native bias.
Figure 4 shows the average performance scores for all models and groups. The figure shows that no clear performance difference exists among the groups compared to the classification results. A slight performance difference favoring the WN group is found for coherence, with the NWN and NN groups performing similarly. Nevertheless, the performance differences are not substantial. Therefore, we conclude that generation tasks are rather robust against (Western) native bias. Nevertheless, when zooming in on the results, we find discrepancies depending on the specific task at hand. These are shown in Appendix L. For two of the datasets, namely Story Cloze and Paraphrase, we find differences in terms of the coherence scores. More specifically, the WN group is here outperforming both the NWN and NN groups. Interestingly, these two tasks also include the smallest written annotations by the prompt writer and generated text by the model. Additionally, when analyzing the CNN DailyMail responses, we find differences in summarization styles among groups. We find how non-native speakers tend to stick closer to the original text when summarizing, while native speakers summarize more freely. Finally, the CODA19 dataset comprises medical articles that utilize specialized medical terminology. Given that most annotators were unfamiliar with this vocabulary, native English speakers (WN and NWN) did not have a specific advantage over non-native speakers. Additionally, research articles are commonly written in English by authors from various backgrounds. Therefore, this specific task might be robust against the native versus non-native preference.
Figure 5: This figure shows the average performance for the different classification tasks per model and group. We see how both GPT models clearly prefer the Western native group, while the other models show similar preference for both native groups for the objective classification task. For the subjective classification tasks, the Western native group is the worst performing group for all models. We adjusted the y-axis to range from 0.65 to 1 for clarity.
(Western) Native bias is model-dependent for the classification tasks.
Figure 5 illustrates that the preference for WN speakers over NWN speakers in objective classification tasks varies by model. Notably, this trend is pronounced in GPT-3.5 and GPT-4o, while Qwen and Claude models show little to no performance difference between WN and NWN speakers. Interestingly, OpenAI’s models even appear to favor NN speakers over NWN speakers. Moreover, it is interesting to see how the Qwen model, developed by Chinese researchers shows almost on par results between both native groups. Additionally, within a model family, the performance disparity increases with model size and overall capability. This aligns with prior research showing a positive correlation between model size and biases, such as gender bias tal-etal-2022-fewer. Furthermore, sclar2023quantifying demonstrate that prompt sensitivity does not decrease as models scale, suggesting that larger models may reinforce rather than mitigate biases. Also for the subjective classification tasks, the results are strongly model-dependent. However, all models do provide the lowest performance for the WN group. For the generative results, on the other hand, all models show similar trends as is shown in Appendix L.
Figure 6: Performance drops when the model is told the prompt writer is non-native rather than native, regardless of the correctness of this information, for objective tasks. Subjective tasks are more robust to this anchoring effect. The figure shows average performance per group and task type based on (in)correct nativeness information averaged over the different models and runs; y-axis is adjusted to 0.65-1 for clarity.
Objective classification tasks are largely affected by adding information about the nativeness of the prompt writer.
Figure 6 shows the effect of providing the model with (in)correct information about the nativeness of the annotator on model performance. This figure clearly shows how the additional information of the nativeness highly affects the results. Adding correct information about the nativeness results in a clear performance preference for the native group, while adding incorrect information results in a preference for the non-native group. Moreover, it not only shows how the performance is influenced by this information, but it also reveals deeply embedded bias towards non-native speakers. Adding this information results in a different performance, where the model focuses more on the initial given information than on the prompt itself. This phenomenon is called anchoring.This term is used for human cognitive bias indicating that a person might insufficiently change its estimates away from an initially provided value (jones2022capturing; tversky1974judgment). This effect is demonstrated in LLMs by jones2022capturing, who found that code generation models modify their outputs to align with related solutions included in the prompt. Moreover, also NGUYEN2024100971 shows how LLM responses are highly influenced by previously given information. Our results reveal a similar anchoring effect, where the model focuses on the additional information about the nativeness of the prompt writer, regardless of whether or not this information is correct. This anchoring effect was most clearly present for Sonnet. We find that Sonnet answered several questions in languages other than English, such as Spanish, French, or Indonesian, when responding as if interacting with non-native speakers. This resulted in a clear drop in performance as is also shown in Figure 11 in Appendix K. Note that this occurred both for native and non-native speakers. From the other models, we see that Qwen and GPT4o seem to be most robust against this added information. GPT3.5 and Haiku did show performance differences, however, not as pronounced as Sonnet. We manually analyzed examples for GPT3.5 and Haiku to gather more insight into the performance difference. GPT3.5 makes more mistakes when informed about the prompt writer being non-native, due to repetition of the instructions, rather than answering the question. Haiku explains the answers, arguing why one option is better than another, thereby failing to follow the instructions. If both answers are mentioned, we classify the response as inaccurate.
Figure 7: Generative tasks are more robust against the anchoring effect. The figure shows average performance per group and metric averaged over the different models and runs based on (in)correct nativeness information; We rescaled the results so that they range from 0 to 1.
The subjective classification tasks and generation tasks are more robust against this additional information about the prompt writer’s nativeness.
In the subjective classification tasks, we observed only slight performance differences, with the non-native group consistently outperforming the native group. These experiments appear largely unaffected by the addition of information, as the non-native group remains the best-performing regardless of whether accurate or inaccurate details about nativeness are introduced. Also for the generative tasks, the addition of information about the prompt writer’s nativeness does not impact performance ranking, as shown in Figure 7. All different groups continue to perform similarly, regardless of the additional information provided.These findings suggest that models are more robust to nativeness cues in generative and subjective classification tasks than in objective classification tasks. This is likely due to their primary optimization for generation rather than classification, particularly given that we use the chat-based versions. Additionally, the longer context in both the initial prompt and generated output may reduce the impact of the anchoring effect.
@@ 6 Discussion
In our experiments, we define native bias as the model’s performance disparity when prompted by native versus non-native English speakers. Additionally, we also further split the native speakers into two groups: Western Natives (WN) and Non-Western Natives (NWN). In general, we find that there are performance differences when the model is prompted by people from different backgrounds.
More specifically, we find an interesting overall preference towards the WN group, where the NWN group is performing similarly as the NN group. This aligns with literature showing how models are tuned towards western native English dialects. The subjective classification tasks on the other hand, favor the western native group the least, across all different models, contradicting findings from santy-etal-2023-nlpositionality; durmus2023towards. This is explained by the models interpreting the (western) native results more positively than intended.
When analyzing the generative results, we find that this task type is more robust and performing similarly for all evaluated groups. This is probably due to the longer context length in both input and generated output, which seems to help the models to perform similarly across different groups. Additionally, the model checkpoints used were also optimized for these generative tasks rather than classification tasks.
We show how the performance ranking of the three groups are also model-dependent for the classification tasks. More specifically, the two GPT models are even preferring the NN group over the NWN group on objective classification tasks. The other models show similar performance for both native groups, for the objective classification tasks. Interestingly, Qwen is thus not showing this clear WN preference, but rather a general native preference, similar to Sonnet and Haiku. This is especially interesting given that Qwen is made by researchers that are not based in the US. In literature, studies showing Western native bias have also been conducted on models made by researchers from Western countries (hofmann2024ai; santy-etal-2023-nlpositionality; durmus2023towards). However, as was shown by buyl2024large, different models have different ideologies, which in turn influence the different biases entailed in the models. Furthermore, also for the subjective tasks, we see how group preference depends on the model. Nevertheless, all models perform worst for the WN group. The generative tasks on the other hand seem to perform similarly across all models.
Finally, we show how a strong anchoring effect occurs when the model is made aware of the nativeness of the prompt writer for the objective classification tasks. The bias is so deeply engraved that informing the models about the nativeness of both groups results in a preference towards the group that was indicated as native, regardless of the correctness of this information, being led by this additional information rather than by the prompt itself. This anchoring effect has been shown to exist in LLMs for a wide range of applications (jones2022capturing; NGUYEN2024100971; echterhoff-etal-2024-cognitive). echterhoff-etal-2024-cognitive analyze the existence of cognitive bias in decision-making with LLMs, while NGUYEN2024100971 focus on using LLMs for financial forecasting. Finally, jones2022capturing focus in a case study on code generation. Our analyses show the existence of this anchoring effect for the objective classification tasks observing differences across models. GPT4o appears most resistant to this anchoring effect, while Sonnet on the other hand even changes the language of the response based on this anchor. echterhoff-etal-2024-cognitive similarly find how GPT4 seems less prone to the anchoring effect than GPT3.5. NGUYEN2024100971 on the other hand, find the opposite for financial forecasting. Nevertheless, given that LLMs are not optimized for this task, this could also affect the conclusions. In our experiments, we also find that the anchoring effect is not clearly present for the generative results, probably due to the optimization of these models towards generative tasks compared to classification tasks, given that we used chat-versions.